# KBAlign: Efficient Self Adaptation on Specific Knowledge Bases

### v3

@@ 4.2 Evaluation Metrics
For LooGLE and ASQA, we consider the evaluation of the original dataset and decide to utilise the following metrics: F1 score, which measures the harmonic mean of precision and recall; Match score, which measures the recall of key elements in long-form answer; BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), which are traditional text generation similarity metrics and evaluate the precision and recall correspondingly; BERT score (Zhang* et al., 2020) calculates cosine similarity to assess semantic consistency, utilizing embeddings generated by the text2vec (Xu, 2023) model from sentences. Additionally, semantic judgment by the representative OpenAI LLM, GPT-4o, is used to evaluate the quality of responses further. Detailed prompts are provided in Section A.
For JEC-QA, we use the official evaluation metrics, where only precise prediction of options could be scored, regardless of whether the questions were single or multiple-choice.
@@ 4.3 Other Settings
For the training process, we tune all parameters of MiniCPM, while conduct a parameter-efficient tuning for LLaMA-3.1, utilizing the LoRA (Hu et al., 2021) strategy to reduce the need for computing power costs. We adopt the mixed-precision training with the BMTrain  [^4] https://github.com/OpenBMB/ModelCenter?tab=readme-ov-file and LLaMA Factory (Zheng et al., 2024) framework to speed up.
For the inference process in both annotation and test, we adopt the bge-large-en-v1.5 model for English materials and bge-base-zh-v1.5 for Chinese (Xiao et al., 2023) as the basic retriever of RAG. In the test scenario, the chunks of KB materials are divided with less than 128 tokens, and the top 8 relevant chunks are provided. To ensure continuity of information, we apply an overlap rate of 15% between consecutive chunks. We adopt vLLM Kwon et al. (2023) to speed up inference.
Methods LooGLE ASQA JEC-QA
F1 BLEU ROUGE BERT LLM Match BLEU ROUGE BERT LLM Single Multi Total
@@
GPT-3.5 35.42 13.11 32.67 80.99 78.08 26.79 3.19 25.37 86.61 51.66 14.49 17.92 16.32
     wo QE 35.27 12.97 32.51 80.94 77.91 27.40 3.13 24.92 86.71 51.52 13.84 19.15 16.68
GPT-4o 40.20 14.93 36.07 81.70 82.93 32.18 2.87 27.69 87.14 67.76 21.95 26.42 24.33
     wo QE 40.21 15.18 36.06 81.71 83.29 32.15 2.94 27.49 87.10 67.88 20.11 27.36 23.98
MiniCPM-2.4B
Original 30.92 10.73 29.27 80.70 64.76 11.91 1.67 21.84 82.30 22.92 39.24 13,87 25.69
     wo QE 30.31 10.50 28.62 80.37 64.72 12.37 1.72 22.36 82.90 22.42 38.38 14.06 25.39
Ours 54.09 18.32 49.75 86.48 75.19 15.68 2.67 24.59 85.41 24.81 49.95 9.94 28.91
     $\Delta$ (+23.17) (+7.59) (+20.48) (+5.78) (+10.43) (+3.77) (+1.00) (+2.75) (+3.11) (+1.89) (+10.71) (-3.93) (+3.22)
     wo QE 53.76 18.55 49.61 86.23 73.19 16.12 2.75 24.65 85.48 25.69 49.41 10.92 29.16
LLaMA3.1-8B-Instruction
Original 40.46 15.91 36.31 81.57 77.15 20.21 2.94 23.24 84.93 37.28 22.70 24.66 23.73
     wo QE 39.94 15.78 36.14 81.50 77.08 20.03 2.94 23.09 85.14 35.64 22.92 24.07 23.53
Ours 62.07 21.73 57.34 88.63 80.16 25.23 3.43 23.59 86.29 42.44 34.59 14.13 23.83
     $\Delta$ (+21.61) (+5.82) (+21.03) (+6.06) (+2.85) (+5.02) (+0.49) (+1.35) (+1.36) (+5.16) (+11.89) (-10.53) (+0.10)
     wo QE 61.79 21.60 57.09 88.55 79.96 25.56 3.23 23.27 86.89 41.31 34.16 14.42 23.78
Table 1: KBAlign adaptation experiments on LooGLE, ASQA, and JEC-QA. We report average for 3 random seeds.
For time cost estimation, we provide the result on ASQA after scaling to the capacity of an A100 GPU: short-dependency annotation for 1k data items takes 30 min, long-dependency annotation for 1k data items takes 140 min, and iterative tuning process takes 160 min.
Hyper-parameters are provided in Section A.
@@ 4.4 Result Analysis
Main Experiments. The main results for our experiments are shown in Table 1. We provide the GPT-series results, the initial version and the self-adapted version of both MiniCPM-2.4B and LLaMA-3.1-8B-Instruction on the three dataset. Overall, comparing the “Ours” lines with the corresponding original model, we can see that KBAlign ensures an obvious improvement on most of the metrics, regardless of the dataset and the backbone model. This shows the generalization and effectiveness of our method.
Nevertheless, our strategies still produce differentiated effects in the three scenarios. For LooGLE which evaluates the master of precise local knowledge, self-annotated tuning brings a huge improvement (over 20% on F1) and the adapted 2.4B model can surpass LLaMA-3.1-8B & GPT-4o performance. For ASQA emphasizing long-form answer that covers global information, however, the improvement is comparably marginal (less than 5% on Match). The first possible reason is that backbone models have already mastered the WikiPedia knowledge in pre-training, and extra adaptation is redundant. The second reason is the over emphasis of local knowledge in the responses, making QE strategy provide even more limited retrieval results. The same trend is also reflected in the single-choice and multiple-choice tests of JEC-QA. Our method easily surpasses some legal-domain models in the former (such as 40.8 single-choice score reported for 7B legal LLM (Wan et al., 2024)), while in the latter the performance even declines slightly due to reasons such as the output format. This indicate the challenge of learning knowledge with a long information span and logical chain.
@@
     wo know 52.16 14.71 47.85 85.75 69.35
     wo RAG 15.40 1.66 15.16 73.54 11.64
     wo QE 53.76 18.55 49.61 86.23 73.19
     wo verify 42.69 15.95 39.57 82.74 72.37
Llama3.1-8B-Instruction
@@
     wo long 13.34 1.18 21.26 81.91 24.32
     wo verify 14.28 2.22 24.32 84.02 23.20
Llama3.1-8B-Instruction
@@
Ablation Study. We assess the effectiveness of our strategies by side experiments in LooGLE and ASQA, and provide the results in Table 2 and 3. To validate the quality of self annotated data, we try to replace the annotation model with GPT-4-turbo (“GPT Data”) in LooGLE, and further replace the data with golden training set (“Golden”) in ASQA. We find that the eventual scores are not much higher than current setting, especially when compared with the original setting without adaptation, proving the usefulness of self annotation.
By conducting ablation study on ASQA, we prove that the long-dependency annotation (“Ours” vs. “wo long”) plays a vital role, in which the comprehensive responses are expected. Considering the higher time cost (about 4 times of short-dependency), we discard this strategy to the local QA task in LooGLE. Meanwhile, self-verify tuning (“Ours” vs. “wo verify”) also helps improve the performance for both dataset by correcting errors of the current stage in a targeted manner.
To explore the mechanism of improvement more clearly, we also conduct a cross-validation on LooGLE (“wo know”), in which the amount of training data keeps the same while the exact information corresponding to the test questions are removed during self-annotation. We can see that the self adaptation still helps refine the performance, but worse than the complete setting. This indicates that domain knowledge from KB and task format is the main reason of the score rising, while the precise information related to the test data also helps the improvement.
@@ 4.5 Case Studies
We display typical cases in Fig. 5 to explain the specific usefulness of our strategies. “Baseline” refers to MiniCPM-2.4B and “ours” refers to the adapted version of it. Overall, KBAlign achieves a general grasp of current KB, a better knowledge answering, and a reasonable confidence verification.
@@ 5 Discussion and Limitations
Still, our approach has some limitations. Firstly, while the current method excels in KBQA tasks, especially those focused on local information within the KB, it offers less support for tasks requiring comprehensive global information analysis. This suggests a need for more refined data annotation strategies. Secondly, training on small-scale targeted data can lead to a reduction in the model’s general domain abilities, such as instruction-following. Mixing specific KB data with general domain data, while helpful, conflicts with our goal of minimizing adaptation time and cost. We may need to explore techniques like model plugins and routing selection to strike a better balance. Lastly, given the strong influence of retrieval quality on QA performance found in our practice, it may be necessary to consider adapting the retriever during specific KB adaptations. Applying self-supervised strategies to retriever training could be a promising direction.
@@ Appendix A Appendix
A.1 Hyper-parameter Settings
For hyper-parameter settings, we conduct a grid search in the vicinity based on the empirical values provided in the sample code of the MiniCPM-2.4B model, and finally determine batch size as $8$ and learning rate as $51e-$ 5. Other settings include warm-up steps as $50$ and weight decay as $0.1$. For the LLaMA-3.1-8B-Instruction model, we adopt a parameter-efficient tuning approach using the LoRA strategy, with alpha as $16$ and rank as $8$. Other settings include a cosine learning rate scheduler with a warm-up ratio of $0.1$ and a weight decay of $0.1$.
A.2 Prompts

### v4

@@ 4.2 Evaluation Metrics
For LooGLE, ASQA and BioASQ, we consider the evaluation of the original dataset and decide to utilise the following metrics: (1) Rule metrics: F1 score, which measures the harmonic mean of precision and recall; Match score, which measures the recall of key elements in long-form answer; For JEC-QA, only precise prediction of options could be scored, regardless of whether the questions were single or multiple-choice.
(2) Similarity metrics: BERT score (Zhang* et al., 2020) calculates cosine similarity to assess semantic consistency, utilizing embeddings generated by the text2vec (Xu, 2023) model from sentences; BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), which are traditional text generation similarity metrics provided in ablation studies.
(3) Intelligent metrics: semantic judgment by the representative OpenAI LLM, GPT-4o, is used to evaluate the quality of responses further. Detailed prompts are provided in Section A.
@@ 4.3 Other Settings
We tune all parameters of MiniCPM, while conduct a parameter-efficient tuning for LLaMA-3.1, utilizing the LoRA (Hu et al., 2021) strategy to reduce the need for computing power costs. In the test scenario, the chunks of KB materials are divided with less than 128 tokens, and the top 8 relevant chunks are provided.
Methods LooGLE ASQA JEC-QA BioASQ
F1 BERT LLM Match BERT LLM Single Multi Total F1 BERT LLM
@@
GPT-3.5 35.42 80.99 78.08 26.79 86.61 51.66 14.49 17.92 16.32 17.80 80.55 93.85
     w/o QE 35.27 80.94 77.91 27.40 86.71 51.52 13.84 19.15 16.68 18.55 80.57 93.23
GPT-4o 40.20 81.70 82.93 32.18 87.14 67.76 21.95 26.42 24.33 31.73 81.31 94.15
     w/o QE 40.21 81.71 83.29 32.15 87.10 67.88 20.11 27.36 23.98 31.39 80.83 94.46
MiniCPM-2B
Vanilla RAG 30.92 80.70 64.76 11.91 82.30 22.92 39.24 13,87 25.69 29.27 82.37 84.92
     w/o QE 30.31 80.37 64.72 12.37 82.90 22.42 38.38 14.06 25.39 30.23 82.71 83.69
RAFT 44.36 84.05 70.73 12.03 85.94 16.18 17.30 14.06 15.57 6.66 81.96 89.23
LM 50.15 84.77 65.62 10.72 81.27 21.03 47.36 7.98 23.73 55.44 88.62 81.85
Ours 54.09 86.48 75.19 15.68 85.41 24.81 49.95 9.94 28.91 61.38 89.95 87.69
     $\Delta$ (+23.17) (+5.78) (+10.43) (+3.77) (+3.11) (+1.89) (+10.71) (-3.93) (+3.22) (+32.11) (+7.58) (+2.77)
     w/o QE 53.76 86.23 73.19 16.12 85.48 25.69 49.41 10.92 29.16 61.91 89.91 89.54
LLaMA3.1-8B-Instruct
Vanilla RAG 40.46 81.57 77.15 20.21 84.93 37.28 22.70 24.66 23.73 27.96 81.55 92.62
     w/o QE 39.94 81.50 77.08 20.03 85.14 35.64 22.92 24.07 23.53 27.74 81.66 91.08
RAFT 42.13 84.88 77.91 23.42 85.92 38.74 23.24 15.47 19.09 44.94 83.36 93.54
LM 54.06 85.53 78.58 19.07 82.52 38.04 20.40 9.57 13.90 56.28 88.02 90.15
Ours 62.07 88.63 80.16 25.23 86.29 42.44 34.59 14.13 23.83 70.97 92.06 93.54
     $\Delta$ (+21.61) (+6.06) (+2.85) (+5.02) (+1.36) (+5.16) (+11.89) (-10.53) (+0.10) (+43.01) (+10.51) (+0.92)
     w/o QE 61.79 88.55 79.96 25.56 86.89 41.31 34.16 14.42 23.78 73.30 92.72 94.48
Table 1: KBAlign adaptation experiments on LooGLE, ASQA, JEC-QA and BioASQ. We report average for 3 random seeds.
Hyper-parameters, retrieval and speed-up settings are provided in Section A.
@@ 4.4 Result Analysis
Time Costs. We first estimate the time cost to prove the efficiency of our method. We provide the result on ASQA after scaling to the capacity of an A100 GPU: short-dependency annotation for 1k data items takes 30 min, long-dependency annotation for 1k data items takes 140 min, and iterative tuning process takes 160 min. Comparably, direct language modeling training takes 480 min, which is longer than the whole process of KBAlign. As for RAFT, it involves larger models and longer CoT responses requiring more annotation time, and the tuning time is controlled to be the same with us.
Main Experiments. Results for our experiments are shown in Table 1. We provide the GPT-series results, the initial version and the self-adapted version of both MiniCPM-2B and LLaMA-3.1-8B-Instruct on the four dataset. Overall, comparing the “Ours” lines with the corresponding vanilla RAG, we can see that KBAlign ensures an obvious improvement on most of the metrics, regardless of the dataset and the backbone model, showing its generalization and effectiveness.
The simple language modeling helps align the model to several KBs and gets marginal promotion, while not always effective, further proving the necessity of self annotation. The advanced baseline RAFT relies on the quality of CoT reasoning, which requires quite large amount of training. When aligned with our high-efficient training setting, its effect is not always obvious.
Task Differences. Nevertheless, our strategies still produce differentiated effects in the four scenarios. For LooGLE which evaluates the master of precise local knowledge, self-annotated tuning brings a huge improvement (over 20% on F1) and the adapted 2B model can surpass LLaMA-3.1-8B & GPT-4o performance. For ASQA emphasizing long-form answer that covers global information, however, the improvement is comparably marginal (less than 5% on Match). The first possible reason is that backbone models have already mastered the WikiPedia knowledge in pre-training, and extra adaptation is redundant. The second reason is the over emphasis of local knowledge in the responses, making QE strategy provide even more limited retrieval results.
The same trend is also reflected in the single-choice and multiple-choice tests of JEC-QA. Our method easily surpasses some legal-domain models in the former (such as 40.8 single-choice score reported for 7B legal LLM (Wan et al., 2024)), while in the latter the performance even declines slightly due to reasons such as the output format. This indicate the challenge of learning knowledge with a long information span and logical chain.
@@
     w/o know 52.16 14.71 47.85 85.75 69.35
     w/o RAG 15.40 1.66 15.16 73.54 11.64
     w/o QE 53.76 18.55 49.61 86.23 73.19
     w/o verify 42.69 15.95 39.57 82.74 72.37
Llama3.1-8B-Instruct
@@
     w/o long 13.34 1.18 21.26 81.91 24.32
     w/o verify 14.28 2.22 24.32 84.02 23.20
Llama3.1-8B-Instruct
@@
Ablation Study. We assess the effectiveness of our strategies by side experiments in LooGLE and ASQA, and provide the results in Table 2 and 3. To validate the quality of self annotated data, we try to replace the annotation model with GPT-4-turbo (“GPT Data”) in LooGLE, and further replace the data with golden training set (“Golden”) in ASQA. We find that the eventual scores are not much higher than current setting, especially when compared with the vanilla setting without adaptation, proving the usefulness of self annotation.
By conducting ablation study on ASQA, we prove that the long-dependency annotation (“Ours” vs. “w/o long”) plays a vital role, in which the comprehensive responses are expected. Considering the higher time cost (about 4 times of short-dependency), we discard this strategy to the local QA task in LooGLE. Meanwhile, self-verify tuning (“Ours” vs. “w/o verify”) also helps improve the performance for both dataset by correcting errors of the current stage in a targeted manner.
To explore the mechanism of improvement more clearly, we also conduct a cross-validation on LooGLE (“w/o know”), in which the amount of training data keeps the same while the exact information corresponding to the test questions are removed during self-annotation. We can see that the self adaptation still helps refine the performance, but worse than the complete setting. This indicates that domain knowledge from KB and task format is the main reason of the score rising, while the precise information related to the test data also helps.
@@ 4.5 Case Studies
We display typical cases in Fig. 5 to explain the specific usefulness of our strategies. “Baseline” refers to MiniCPM-2B and “ours” refers to the adapted version of it. Overall, KBAlign achieves a general grasp of current KB, a better knowledge answering, and a reasonable confidence verification.
@@ 5 Discussion and Limitations
Still, our approach has some limitations: (1) Global information: While the current method excels in KBQA tasks, especially those focused on local information within the KB, it offers less support for tasks requiring comprehensive global information analysis. This suggests a need for more refined data annotation strategies.
(2) General Capability: Training on small-scale targeted data can lead to a reduction in the model’s general domain abilities, such as instruction-following. Mixing specific KB data with general domain data, in fact, has proved to be helpful in our side experiment, which is displayed in section A. However, this conflicts with our goal of minimizing adaptation time and cost. We may need to explore techniques like model plugins and routing selection to strike a better balance.
(3) Retriever Adaptation: Given the strong influence of retrieval quality on QA performance found in our practice, it may be necessary to consider adapting the retriever during specific KB adaptations. Applying self-supervised strategies to retriever training could be a promising direction.
@@ Appendix A Appendix
A.1 Dataset Details
LooGLE (Li et al., 2023). We use the short-dependency data in LooGLE for retrofitting, combining altogether 2.2M tokens of text as $K$, and the corresponding 1,951 Q&A pairs for test.
ASQA (Stelmakh et al., 2022). For each question, there exists several related segments of text from WikiPedia, and a comprehensive long answer that covers much information from them. We collect 794 Q&A pairs for test, their targeted segments and other related passages from WikiPedia, and get 1.8M tokens of text as $K$.
JEC-QA (Zhong et al., 2020). Related laws and reference books are seen as $K$, including 21M tokens of text. The train set has also not been used for adaptation, and the test set contains 1,985 of multiple choices.
BioASQ  [^5] https://huggingface.co/datasets/kroshan/BioASQ. This contains 324 English questions for testing, each annotated with relevant documents, snippets, and both exact and ideal answers. We utilize 0.41M tokens of text as $K$.
A.2 Detailed Settings
For hyper-parameter settings, we conduct a grid search in the vicinity based on the empirical values provided in the sample code of the MiniCPM-2B model, and finally determine batch size as $8$ and learning rate as $51e-$ 5. Other settings include warm-up steps as $50$ and weight decay as $0.1$. For the LLaMA-3.1-8B-Instruct model, we adopt a parameter-efficient tuning approach using the LoRA strategy, with alpha as $16$ and rank as $8$. Other settings include a cosine learning rate scheduler with a warm-up ratio of $0.1$ and a weight decay of $0.1$.
For the training process, we adopt the mixed-precision training with the BMTrain  [^6] https://github.com/OpenBMB/ModelCenter?tab=readme-ov-file and LLaMA Factory (Zheng et al., 2024) framework to speed up.
For the inference process in both annotation and test, we adopt the bge-large-en-v1.5 model for English materials and bge-base-zh-v1.5 for Chinese (Xiao et al., 2024) as the basic retriever of RAG. To ensure continuity of information, we apply an overlap rate of 15% between consecutive chunks. We adopt vLLM (Kwon et al., 2023) to speed up inference.
A.3 Prompts
@@ A.4 Supplementary Case
A.5 General Domain Performance
We conduct experiments on the widely adopted MMLU benchmark (Hendrycks et al., 2021) to evaluate the general ability reduction of KBAlign brought to the backbone model. w mix refers to mixing general instruction tuning data (Ding et al., 2023) with specific KB data. As shown in Table 4, tuning on mixed data could also achieve most of the downstream improvement with a fairly small general performance degradation.
Methods MMLU LooGLE F1
MiniCPM-2B
Vanilla RAG 44.07 30.92
Ours 38.54 54.09
     w mix 43.45 52.84
Llama3.1-8B-Instruct
Vanilla RAG 57.38 40.46
Ours 49.41 62.07
     w mix 54.61 61.33
Table 4: Balance between domain knowledge and general capacity.
Algorithm 1 KBAlign Framework
1:Model $M$, Retriever $R$, Golden context $C_{g}$, Question $Q$, Answer $A$, Split size $k$
2:Fine-tuned model $M_{k}$
3:Annotation Process:
4:procedure ShortAnnotation($C_{g}$)
5:     $Q_{\mathrm{short}}\leftarrow M(C_{g})$
6:     $C_{R}\leftarrow R(Q_{\mathrm{short}})$
7:     $C\leftarrow C_{g}\oplus C_{R}$
8:     $A_{\mathrm{short}}\leftarrow M(Q_{\mathrm{short}}\oplus C)$
9:     return $\langle Q_{\mathrm{short}},A_{\mathrm{short}}\rangle$
10:end procedure
11:procedure LongAnnotation($\{S_{i}\}_{i=1}^{n}$)
12:     $C_{g}\leftarrow\bigoplus_{i=1}^{n}S_{i}$
13:     $Q_{\mathrm{long}}\leftarrow M(C_{g})$
14:     $C_{R}\leftarrow R(Q_{\mathrm{long}})$
15:     for $i=1,\ldots,n$ do
16:         $C_{i}\leftarrow S_{i}\oplus C_{R}$
17:         $I_{i}\leftarrow M(Q_{\mathrm{long}}\oplus C_{i})$
18:     end for
19:     $A_{\mathrm{long}}\leftarrow M(Q_{\mathrm{long}}\oplus\bigoplus_{i=1}^{n}I_{i})$
20:     return $\langle Q_{\mathrm{long}},A_{\mathrm{long}}\rangle$
21:end procedure
22:Training Phase:
23:Split annotated data $\{\langle Q,A\rangle\}$ into $k$ parts $\{\langle Q_{i},A_{i}\rangle\}_{i=1}^{k}$
24:Initial Tuning:
25:$\mathcal{L}_{1}=0.5\mathbb{E}[\|M(Q_{1})-A_{1}\|]+0.5\mathbb{E}[\|M(Q_{1}\oplus R(Q_{1}))-A_{1}\|]$
26:$M_{1}\leftarrow\arg\min_{M}\mathcal{L}_{1}$
27:Iterative Verifying:
28:for $i=2$ to $k$ do
29:     $C_{R}\leftarrow R(Q_{i})$
30:     $P_{i}\leftarrow M_{i-1}(Q_{i}\oplus C_{R})$
31:     $V_{i}\leftarrow M_{i-1}(Q_{i}\oplus P_{i}\oplus A_{i})$
32:     $\mathcal{L}_{i}=0.375\mathbb{E}[\|M(Q_{i})-A_{i}\|]+0.375\mathbb{E}[\|M(Q_{i}\oplus C_{R})-A_{i}\|]$
33:     $\quad+0.125\mathbb{E}[\|M(Q_{i}\oplus P_{i})-V_{i}\|]+0.125\mathbb{E}[\|M(Q_{i}\oplus C_{R}\oplus P_{i})-V_{i}\|]$
34:     $M_{i}\leftarrow\arg\min_{M_{i-1}}\mathcal{L}_{i}$
35:end for